Closed Bug 848992 Opened 7 years ago Closed 7 years ago

Re-image r4 Mac foopys as Lion (10.7) test slaves

Categories

(Infrastructure & Operations :: DCOps, task)

x86_64
macOS
task
Not set

Tracking

(Not tracked)

RESOLVED FIXED

People

(Reporter: coop, Unassigned)

References

Details

This is currently blocked on Linux foopy replacements that will be deployed as part as bug 666044 (or dependent bugs).

Given that this will be blocked for a while, I'm not assigning it to relops/dcops yet.

The machines in question are:
foopy12
foopy13
foopy14
foopy15
foopy16
foopy17
Assignee: nobody → server-ops-releng
Component: Release Engineering: Platform Support → Server Operations: RelEng
QA Contact: coop → arich
This doesn't depend on 666044 at all unless you're not going to use the linux foopies before then.  It should depended on 851579 or, more appropriately, whatever you guys have on the releng to correspond to setting stuff up after that bug was closed.
(In reply to Amy Rich [:arich] [:arr] from comment #1)
> This doesn't depend on 666044 at all unless you're not going to use the
> linux foopies before then.  It should depended on 851579 or, more
> appropriately, whatever you guys have on the releng to correspond to setting
> stuff up after that bug was closed.

Sorry, underlying assumptions have changed since this was filed. Removing dependency.
No longer depends on: 666044
Depends on: 851579
Depends on: 850317
dcops, please relocate these machines (new names listed below) to scl1 and change the rack and switch information in inventory and add a PDU k/v store for whatever IP addressable PDU these are hooked up to.  Please also netboot them so that they reinstall with the proper OS.


foopy12 -> talos-r4-lion-085
foopy13 -> talos-r4-lion-086
foopy14 -> talos-r4-lion-087
foopy15 -> talos-r4-lion-088
foopy16 -> talos-r4-lion-089
foopy17 -> talos-r4-lion-090
Assignee: server-ops-releng → server-ops-dcops
colo-trip: --- → mtv1
Component: Server Operations: RelEng → Server Operations: DCOps
QA Contact: arich → dmoore
Note: these also need to be hooked up to EDID.
Depends on: 721386
Hosts have been unracked at MV office.  Transporting to SCL1.
colo-trip: mtv1 → scl1
All 6 hosts have been racked and reimaged.  We only have two available EDIDs to hook up at the moment.

EDID enabled:
talos-r4-lion-089
talos-r4-lion-090


Awaiting for EDID shipment:
talos-r4-lion-085
talos-r4-lion-086
talos-r4-lion-087
talos-r4-lion-088
EDID adapters have shipped and should arrive on 4/4:

UPS 1ZRV25160348814224
UPS 1ZRV25160352231239
All hosts are now connected to EDIDs.
Status: NEW → RESOLVED
Closed: 7 years ago
Resolution: --- → FIXED
Product: mozilla.org → Infrastructure & Operations
You need to log in before you can comment on or make changes to this bug.